Find the smallest positive inverse of 10 (mod 17)
Answer #1
10* 0  % 17 = 0
10* 1  % 17 = 10
10* 2  % 17 = 3
10* 3  % 17 = 13
10* 4  % 17 = 6
10* 5  % 17 = 16
10* 6  % 17 = 9
10* 7  % 17 = 2
10* 8  % 17 = 12
10* 9  % 17 = 5
10* 10  % 17 = 15
10* 11  % 17 = 8
10* 12  % 17 = 1


So, the answer is 12

An swer

12

\color{red}Please\; upvote\;the \;solution \;if \;it \;helped.\;Thanks!